OpenStudy (anonymous):

is this correct? sqrt20a^2b=2asqrt5b

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you. just wanted to make sure!

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

If the problem expression is the following: \[\sqrt{20a^{2b}}=2a\sqrt{5b} \] then the lhs is not equal to the rhs because \[2 \sqrt{5} a^b \neq 2 \sqrt{5} a \sqrt{b} \]

OpenStudy (anonymous):

thats not the expression

OpenStudy (anonymous):

What is you understanding regarding the form of the original?

OpenStudy (anonymous):

it is \[\sqrt{20a ^{2}b}\]

OpenStudy (anonymous):

Thank you for responding.